OpenStudy (vshiroky):

decide if the ordered pair (3,-2) is a solution to the equation 2x-y=4. yes or no?

jimthompson5910 (jim_thompson5910):

Replace x with 3 and y with -2. This is because all points are of the form (x,y) 2x-y=4 2(3)-(-2)=4 6 + 2 = 4 8 = 4 ... this is FALSE, so (3,-2) is NOT a solution to 2x-y=4 In other words, the point (3,-2) is NOT on the line 2x-y=4

OpenStudy (vshiroky):

Thanks Jim. Thank is what I thought, but for some reason I always get confused to change the 2nd to a positive or not so I wanted to make sure haha.

jimthompson5910 (jim_thompson5910):

Just remember that negatives are like saying opposites So -(-2) is the same as "the opposite of -2" which is really +2 Another way is to think of the two negatives cancelling to make a positive